# Notions of entropy for unitary representations

### v2

@@ Abstract
Several notions of entropy are studied widely in the ergodic theory of probability-preserving actions of countable groups. These include the generalization of Kolmogorov–Sinai entropy for amenable groups using Følner sequences, percolative entropy entropy for general countable groups, and Bowen’s sofic entropy for sofic groups.
In this work we pursue these notions across the analogy between ergodic theory and representation theory. We arrive at new quantities associated to unitary representations of groups and representations of other C*-algebras. Our main results show that these new quantities can generally be evaluated as Fuglede–Kadison determinants. The resulting formulas for those determinants can be seen as various non-commutative generalizations of Szegő’s limit theorem for Toeplitz determinants. They also establish connections with Arveson’s theory of subdiagonal subalgebras, and with some more recent exact entropy calculations in the ergodic theory of actions by automorphisms of compact Abelian groups.
@@ 1.1 A Szegő limit theorem over amenable groups
If $\varphi$ is a positive definite function on $\mathbf{Z}$, then Bochner’s theorem identifies it as the Fourier–Stieltjes transform of a finite Borel measure $\mu$ on the circle group $\mathbf{T}$. Let $m_{\mathbf{T}}$ be the Lebesgue probability measure on $\mathbf{T}$. In this context, Szegő’s limit theorem identifies the asymptotic behaviour of finite Toeplitz determinants approximating $\varphi$:
@@
Let $D_{n}$ be the determinant of the Toeplitz matrix $[\varphi(i-j)]_{i,j=1}^{n}$ for each $n$, and let $\mu_{\mathrm{ac}}$ is the absolutely continuous part of $\mu$. Then
@@
taking the right-hand side to be $\exp(-\infty):=0$ if necessary. ∎
A thorough account of Theorem S and many variations can be found in standard texts such as [Sim05a, Sim05b], which I generally follow for terminology in the sequel. See [Sim05a, Chapter 1] for a broader overview of this area and a history for some of the key contributions. In fact, Szegő’s early works contain (1.1) only when $\mu\ll m$, with the extension to general measures $\mu$ accomplished first by Verblunsky. Theorem S also has a generalization in which $\mu$ and $\varphi$ take values among positive definite $k$-by-$k$ matrices [Sim05a, Section 2.13].
The right-hand side of (1.1) admits the following ‘non-commutative’ point of view. Let $f:=d\mu_{\mathrm{ac}}/dm_{\mathbf{T}}$, so this is an element of $L^{1}(\mathbf{T})$. We can regard $L^{\infty}(\mathbf{T})$ as a von Neumann algebra of multiplication operators on $L^{2}(\mathbf{T})$, and integration with respect to $m_{\mathbf{T}}$ defines a tracial state on this von Neumann algebra. If $f$ is essentially bounded, then the multiplication operator $M_{f}$ belongs to this von Neumann algebra, and more generally $M_{f}$ can be defined as an operator affiliated to that algebra. Then the right-hand side of (1.1) is the logarithm of the Fuglede–Kadison determinant of $M_{f}$ with respect to $m_{\mathbf{T}}$ (see [FK52] or [Dix81, Section I.6.11]). We review these concepts more carefully in Section 3 below. In the context of Szegő’s theorem, this observation goes back at least to [Arv67], where Arveson proposed a non-commutative generalization to ‘subdiagonal subalgebras’, which are discussed further below. It also appears in [Den06, Sectoin 2].
Generalizations of Theorem S to positive definite functions on $\mathbf{Z}^{d}$ satisfying various additional hypotheses are proved in [HL58, Lin75, Dok84]. More recently, further generalizations appear in works by Deninger [Den06, Den09, Den12], Li [Li12] and Li and Thom [LT14] for groups that are amenable but not necessarily Abelian, again with additional hypotheses on the positive definite functions.
Our first main result is a version of this theorem for arbitrary matrix-valued positive definite maps on any countable amenable group $\Gamma$ (see Section 5 for the definition of these maps).
One of the many equivalent characterizations of amenability is that $\Gamma$ has a sequence $(F_{n})_{n\geq 1}$ of finite subsets satisfying the right Følner condition:
@@
The analogous left Følner condition is explained in [Gre69, Section 3.6]. The inversion map of $\Gamma$ converts one condition to the other, so both are characterizations of amenability. We refer to a sequence of subsets that satisfies (1.2) as a right Følner sequence.
Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, and for each finite subset $F$ of $\Gamma$ consider the $F$-by-$F$ block matrix
@@
If $\varphi$ is associated to a representation $\pi$ by the vectors $v_{1}$, $v_{2}$, …, $v_{k}$, then $\varphi[F]$ is the Gram matrix of the tuple of vectors
@@
Let $\lambda$ be the left regular representation. Classic results from functional analysis give a unique decomposition of $\varphi$ as $\varphi_{\mathrm{sing}}+\varphi_{\mathrm{ac}}$, where the GNS representation of $\varphi_{\mathrm{sing}}$ is disjoint from $\lambda$ and the GNS representation of $\varphi_{\mathrm{ac}}$ is a subrepresentation of $\lambda$ (see Propositions 3.3 and 4.6 below). In addition, the part $\varphi_{\mathrm{ac}}$ can be represented in terms of a self-adjoint operator affiliated to $\lambda(\Gamma)^{\prime}$ (Propositions 3.5 and 4.7), and using this we can define the Fuglede–Kadison determinant $\Delta\varphi_{\mathrm{ac}}$ (Definition 4.9).
@@
The literature contains a number of precedents lying between Szegő’s theorem and Theorem A. For example, [Den06, Theorem 3.2] implies the case of Theorem A in which $\varphi_{\mathrm{sing}}=0$ and $\varphi$ can be expressed in terms of $\tau$ and a positive invertible element of the von Neumann algebra $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$. We discuss the previous literature more fully in Subsection 6.3.
@@ 1.2 Analogies with entropy and ergodic theory
More formally, log-determinants arise as values of differential rather than discrete entropy. If $(X_{1},\dots,X_{k})$ is a multivariate Gaussian random vector with covariance matrix $Q$, then its differential Shannon entropy equals $\log\det Q$ up to normalization [CT06, Theorem 8.4.1]. From this point of view, the analogy sketched above is really between discrete and differential entropy. However, the present paper concerns a purely ‘linear’ setting, so we generally discuss this analogy without introducing the extraneous construct of a Gaussian random vector. Either way, a few key differences must be kept in mind when comparing with discrete entropies, particularly these: (i) log-determinants (differential entropies) depend on the ambient dimension in which a (random) vector takes it values; (ii) log-determinants (differential entropies) can have either sign, or even equal $\infty$- ∞.
@@
Now fix a countable group $\Gamma$, and consider actions of two different kinds: measure-preserving actions on probability spaces; and unitary actions on Hilbert spaces. These two settings enjoy a $\Gamma$-equivariant version of the analogy between Gram matrices and joint distributions. Observables (equivalently, finite measurable partitions) on a probability space are analogous to vectors in a Hilbert space, or more generally to finite tuples of vectors. If $(X,\mu,T)$ is a measure-preserving $\Gamma$-system and $\alpha:X\to A$ is an observable, then it generates the shift-system $(A^{\Gamma},\alpha^{\Gamma}_{\ast}\mu,S)$, where $\alpha^{\Gamma}_{\ast}\mu$ is the law of the $\Gamma$-indexed stochastic process $(\alpha\circ T^{g}:\ g\in\Gamma)$. Analogously to this, if $\pi$ is a unitary representation of $\Gamma$ and $v\in H_{\pi}$, then these define the positive definite function $\langle\pi(\cdot)v,v\rangle$ on $\Gamma$, and this construction can be extended to finite tuples of vectors by considering matrix-valued positive definite functions.
@@
This analogy has long stimulated research in both areas. For example, it is clear in Kechris’ adaptation of the relation of weak containment to measure-preserving systems [Kec10, BK20]. Within it, the logarithm of the limit on the left-hand side of (1.3) is the analog of the definition of Kolmogorov–Sinai entropy for a stationary process over an amenable group [Kie75]. Theorems B and C below concern the analogs of two other notions of entropy from ergodic theory.
@@
The analogy between ergodic theory and representation theory crystallizes into at least two different formal relationships. On the one hand, any measure-preserving system gives rise to its Koopman representation [Kec10, Section 10]. On the other hand, any orthogonal real representation of $\Gamma$ can be used to construct a measure-preserving action on a Gaussian Hilbert space, and this construction can be adapted to start with a unitary representation instead [Kec10, Appendices C–E]. However, the first of these relationships does not correctly connect the notions of entropy that we study in this work, and the second introduces unnecessary complications.
@@
On the other hand, if we start with a unitary representation and construct the associated Gaussian system, then there are cases in which a suitable limit of log-determinants for the former should equal a ‘differential’ analog of Kolmogorov–Sinai entropy for the latter. For single transformations or actions of $\mathbf{Z}^{d}$, ‘differential’ Kolmogorov–Sinai entropy for stationary $\mathbf{R}$-valued processes is studied in [HPT08], prompted by earlier work in information theory such as [Pin64]. Section 6 of [HPT08] includes some exact calculations for Gaussian systems and linear transformations between them which boil down to applications of Theorem S itself. However, this differential version of Kolmogorov–Sinai entropy starts to behave quite wildly beyond those Gaussian examples, and for those examples alone we might as well stay within the setting of representation theory. Overall, an analogy at the level of intuition seems more revealing than either the Koopman or the Gaussian construction for our work below.
Beyond Theorem A, we consider to two other notions of entropy from ergodic theory, and develop their analogs for unitary representations of groups or representations of other C*-algebras. The second of these notions is Kieffer’s definition of entropy using a ‘random past’ from [Kie75, Theorem 3], which leads to Theorem B. The third is Bowen’s ‘sofic entropy’ from [Bow10b], which leads to Theorem C. In both cases, we show that these notions of entropy are given by $\log\Delta\varphi_{\mathrm{ac}}$ for some positive functional $\varphi$, except in Theorem C a certain degeneracy may occur and then the entropy equals $\infty$- ∞ (also reflecting a possible behaviour of sofic entropy).
Each of these theorems may be seen as a different ‘non-commutative’ version of Theorem S. In ergodic theory, various predecessors of our results are already discussed this way in the literature, for example in [Lyo05, LT14, Hay16].
@@ 1.3 Subdiagonal subalgebras and random orders
where $P_{n}$ is the orthogonal projection from $L^{2}(\mu)$ to $\mathrm{span}\{z,\dots,z^{n}\}$ [Sim05a, Theorem 1.5.11]. It follows that
@@
This orthogonal projection is the closest point of $N$ to the function $1_{\mathbf{T}}$, so Szegő’s limit theorem is equivalent to
@@
This is the form in which Szegő’s theorem most often appears in the literature on analytic functions, such as in [Gam69, Section V.8] or [Gar07, Theorem IV.3.1]. Those references give essentially Szegő’s own proof, which he published first for the special case $\mu\ll m$ and then much later for the general case by incorporating arguments of Kolmogorov and Krein to handle $\mu_{\mathrm{sing}}$. Two further variations on this proof are recounted in [Sim05a, Sections 2.4 and 2.5]. Meanwhile, Verblunsky gave the first proof of the general case of Theorem S in [Ver35b, Ver35a] (see also [Sim05a, Section 2.3]).
@@
See, for instance, [Wal82, Theorem 4.14].
To generalize Szegő’s theorem and related results to positive definite functions on $\mathbf{Z}^{d}$, one can use the ‘past’ defined by a lexicographic ordering. This approach is developed by Helson and Lowdenslager in [HL58, HL61] (with some later refinements in [Lyo03]). See also [Lin75, Dok84] for alternative proofs under some additional assumptions on the positive definite functions. In ergodic theory, the same use of the lexicographic ordering appears in some early works on the entropy of measure-preserving $\mathbf{Z}^{d}$-actions, such as [Con73, KW72].
This idea has a further natural generalization to any countable group $\Gamma$ that admits a left-invariant total order. But some countable groups do not, including all groups that are not torsion-free. To remove the need for this assumption, one can instead couple to a stationary random ordering of the group. All countable groups admit at least one of these: the ‘Bernoulli random order’ described in Example 7.5.
In ergodic theory, conditioning on the ‘past’ of a stationary random order serves as a replacement for the formula (1.5). If $\Gamma$ is amenable, this idea gives rise to another classic formula for the entropy rate of a process due to Kieffer [Kie75], but the situation is more complicated for non-amenable groups. Our next main theorem develops the analog of this idea for positive definite functions. Here we find once again that the resulting quantity always agrees with the expected Fuglede–Kadison determinant, even if $\Gamma$ is not amenable.
Before formulating the theorem precisely, let us motivate it further via a related finite-dimensional calculation. Let $n$ and $k$ be positive integers. Consider a tuple of $nk$ vectors in some Hilbert space $H$ indexed as an $n$-by-$k$ array, say
@@
Formula (1.6) remains valid under any re-ordering of the tuples $V_{1}$, …, $V_{n}$, and so we can take an expectation over a uniform random order on the right-hand side. Then the symmetry of the random order lets us replace the average of $n$ terms on the right-hand side with a single average over orders. Using $\omega$ to denote a permutation of $1$, …, $n$, and now writing $R_{\omega}$ for the orthogonal projection onto the random subspace
@@
From the viewpoint of ergodic theory, Theorem B is somewhat surprising, because for a non-amenable group $\Gamma$ it gives much better behaviour than the original notion of random-order entropy for a finite-valued stationary processes. Random-order entropy does agree with Rokhlin or sofic entropy (discussed below) for special examples of processes, such as those in [AP18, Alp]. But in general it is only an upper bound for those quantities, and unlike them it need not be invariant under isomorphism [Sew, Section 7]. However, for a positive definite function $\varphi$ on a group $\Gamma$ for which both Theorems B above and C below can be applied, they always give the same value $\log\Delta\varphi_{\mathrm{ac}}$ for their respective notions of entropy. Since Theorem B concerns the analog of percolative entropy, and Theorem C concerns the analog of sofic entropy (explained below), this suggests that the worst pathologies from ergodic theory do not have representation theoretic versions.
It turns out that most of the work towards Theorem B can already be found in the literature. It takes place in the context of Arveson’s theory of ‘subdiagonal subalgebras’ of finite von Neumann algebras [Arv67]. These are a noncommutative generalization of the inclusion of the Hardy space $H^{\infty}(\mathbf{T})$ in $L^{\infty}(\mathbf{T})$. As such, they provide an abstract notion of a ‘past’ in a general finite von Neumann algebra. We recall their definition precisely in Subsection 7.5.
Arveson conjectures a generalization of Szegő’s theorem to this noncommutative setting in [Arv67, item 4.4($\gamma$)]. As far as I know, this is the first reference to suggest that the log-integral should be replaced with a Fuglede–Kadison determinant. Arveson proves this generalization in a few of his examples in [Arv67, Section 5], including some that recover Helson and Lowdenslager’s results for $\mathbf{Z}^{d}$. He also shows how to construct subdiagonal subalgebras for any countable group with an invariant total order, but does not prove his Szegő formula in this generality.
Arveson’s conjectured Szegő theorem was finally proved in general by Labuschagne in [Lab05], in the equivalent form of an abstract ‘Jensen formula’. We recall this theorem as Theorem 7.3 below.

### v3

# Entropy and determinants for unitary representations

@@ Abstract
Several notions of entropy have been studied widely in the ergodic theory of probability-preserving actions of countable groups. These include the generalization of Kolmogorov–Sinai entropy to amenable groups based on Følner sequences, percolative entropy for general countable groups, and Bowen’s sofic entropy for sofic groups.
In this work we pursue these notions across an analogy between ergodic theory and representation theory. We arrive at new quantities associated to unitary representations of groups and representations of other C*-algebras. Our main results show that these new quantities can often be evaluated as Fuglede–Kadison determinants. The resulting determinantal formulas offer various non-commutative generalizations of Szegő’s limit theorem for Toeplitz determinants. They make contact with Arveson’s theory of subdiagonal subalgebras, and also with some exact entropy calculations in the ergodic theory of actions by automorphisms of compact Abelian groups.
[…]
0.4pt0cm
@@ 1.1 A Szegő limit theorem over amenable groups
If $\varphi$ is a positive definite function on $\mathbf{Z}$, then Bochner’s theorem identifies it as the Fourier–Stieltjes transform of a finite Borel measure $\mu$ on the circle group $\mathbf{T}$. Let $m_{\mathbf{T}}$ be the Lebesgue probability measure on $\mathbf{T}$. In this context, Szegő’s limit theorem describes the asymptotic behaviour of finite-dimensional Toeplitz determinants obtained from $\varphi$:
@@
Let $D_{n}$ be the determinant of the Toeplitz matrix $[\varphi(i-j)]_{i,j=1}^{n}$ for each $n$, and let $\mu_{\mathrm{ac}}$ be the absolutely continuous part of $\mu$. Then
@@
taking the right-hand side to be $\exp(-\infty)=0$ if necessary. ∎
A thorough account of Theorem S can be found in standard texts such as [136, 137], which I generally follow for terminology in the sequel. See [136, Chapter 1] for a broader overview of this area and a history for some of the key contributions. In fact, Szegő’s early works contain (1.1) only when $\mu\ll m$, with the extension to general measures $\mu$ accomplished first by Verblunsky. Theorem S also has a generalization in which $\mu$ and $\varphi$ take values among positive definite $k$-by-$k$ matrices [136, Section 2.13].
The right-hand side of (1.1) admits the following ‘non-commutative’ point of view. Let $f:=d\mu_{\mathrm{ac}}/dm_{\mathbf{T}}$, so this is an element of $L^{1}(\mathbf{T})$. We can regard $L^{\infty}(\mathbf{T})$ as a von Neumann algebra of multiplication operators on $L^{2}(\mathbf{T})$, and then integration with respect to $m_{\mathbf{T}}$ defines a tracial state on $L^{\infty}(\mathbf{T})$. If $f$ is essentially bounded, then the multiplication operator $M_{f}$ belongs to this von Neumann algebra; in general, $M_{f}$ can be defined as an operator affiliated to that algebra. Then the right-hand side of (1.1) is the logarithm of the Fuglede–Kadison determinant of $M_{f}$ with respect to $m_{\mathbf{T}}$. We review these concepts more carefully in Section 3 below. In the context of Szegő’s theorem, this observation goes back at least to [9], where Arveson proposed a non-commutative generalization to ‘subdiagonal subalgebras’, which are discussed further below.
Under various additional hypotheses, generalizations of Theorem S to positive definite functions on $\mathbf{Z}^{d}$ were obtained in [85, 105, 51]. More recently, further generalizations appeared in works by Deninger [43, 44, 45], Li [103] and Li and Thom [104]. They handled more general amenable groups, but still with additional hypotheses on the positive definite functions.
For a countable group $\Gamma$, one of the many equivalent conditions for amenability is the existence of a sequence $(F_{n})_{n\geq 1}$ of finite subsets satisfying the right Følner condition:
@@
This condition is discussed in [94, Section II.5], or the analogous left Følner condition is explained in [75, Section 3.6]. The inversion map on $\Gamma$ converts one condition to the other, so both are characterizations of amenability. We refer to a sequence of subsets that satisfies (1.2) as a right Følner sequence.
Our first main result is a version of Theorem S for arbitrary matrix-valued positive definite maps on a countable amenable group $\Gamma$ (see Section 5 for the definition of these maps). Let $\varphi:\Gamma\to\mathrm{M}_{k}$ be positive definite, and for each finite subset $F$ of $\Gamma$ consider the $F$-by-$F$ block matrix
@@
If $\varphi$ is associated to a unitary representation $\pi$ by the vectors $v_{1}$, $v_{2}$, …, $v_{k}$, then $\varphi[F]$ is the Gram matrix of the tuple of vectors
@@
Let $\lambda$ be the left regular representation. Classic results from functional analysis give a unique decomposition of $\varphi$ as $\varphi_{\mathrm{sing}}+\varphi_{\mathrm{ac}}$, where the minimal dilation of $\varphi_{\mathrm{sing}}$ is disjoint from $\lambda$ and the minimal dilation of $\varphi_{\mathrm{ac}}$ is a subrepresentation of $\lambda$ (see Propositions 3.3 and 4.9 below). In addition, $\varphi_{\mathrm{ac}}$ can be represented in terms of a self-adjoint operator affiliated to $\lambda^{\oplus k}(\Gamma)^{\prime}$ (see Propositions 3.5 and 4.10), and using this we can define the Fuglede–Kadison determinant $\Delta\varphi_{\mathrm{ac}}$ (see Definition 4.12).
@@
The literature contains a number of precedents lying between Szegő’s theorem and Theorem A. For example, [43, Theorem 3.2] implies the case of Theorem A in which $\varphi_{\mathrm{sing}}=0$ and $\varphi$ can be expressed in terms of a positive invertible element of the von Neumann algebra $\mathrm{M}_{k}(\lambda(\Gamma)^{\prime\prime})$. We discuss the previous literature more fully in Subsection 6.3.
@@ 1.2 Analogies with entropy and ergodic theory
More formally, log-determinants arise as values of differential rather than discrete entropy. If $(X_{1},\dots,X_{k})$ is a multivariate Gaussian random vector with covariance matrix $Q$, then its differential Shannon entropy equals $\log\det Q$ up to a normalization [40, Theorem 8.4.1]. From this point of view, the analogy sketched above is really between discrete and differential entropy. However, the present paper concerns a purely ‘linear’ setting, so we generally discuss this analogy without the extraneous construct of a Gaussian random vector. Either way, a few key differences must be kept in mind when comparing with discrete entropies, particularly these: (i) log-determinants (differential entropies) depend on the ambient dimension in which a (random) vector takes it values; (ii) log-determinants (differential entropies) can have either sign, or even equal $\infty$- ∞.
@@
Now let $\Gamma$ be a countable group, and consider actions of two different kinds: measure-preserving actions on probability spaces, and unitary actions on Hilbert spaces. These two settings enjoy a $\Gamma$-equivariant version of the analogy between joint distributions and Gram matrices. Observables (equivalently, finite measurable partitions) on a probability space are analogous to vectors in a Hilbert space, or more generally to finite tuples of vectors. If $(X,\mu,T)$ is a measure-preserving $\Gamma$-system and $\alpha:X\to A$ is an observable, then it generates the shift-system $(A^{\Gamma},\alpha^{\Gamma}_{\ast}\mu,S)$, where $\alpha^{\Gamma}_{\ast}\mu$ is the law of the $\Gamma$-indexed stochastic process $(\alpha\circ T^{g}:\ g\in\Gamma)$. Analogously to this, if $\pi$ is a unitary representation of $\Gamma$ and $v\in H_{\pi}$, then these define the positive definite function $\langle\pi(\cdot)v,v\rangle$ on $\Gamma$, and this construction can be extended to finite tuples of vectors by considering matrix-valued positive definite functions.
@@
This analogy has long stimulated research in both areas. For example, it underlies Kechris’ adaptation of the relation of weak containment to measure-preserving systems [93, 31]. Within this analogy, the logarithm of the limit on the left-hand side in Theorem A is the analog of the Kolmogorov–Sinai entropy of a stationary process over an amenable group [96].
@@
The analogy between ergodic theory and representation theory crystallizes into at least two different formal relationships. On the one hand, any measure-preserving system gives rise to its Koopman representation [93, Section 10]. On the other hand, any orthogonal real representation of $\Gamma$ can be used to construct a measure-preserving action on a Gaussian Hilbert space, and this construction can be adapted to start with a unitary representation instead [93, Appendices C–E]. However, the first of these relationships does not correctly connect the notions of entropy that we study in this work, and the second introduces unnecessary complications.
@@
On the other hand, if we start with a unitary representation and construct the associated Gaussian system, then there are cases in which a suitable limit of log-determinants for the former should equal a ‘differential’ analog of Kolmogorov–Sinai entropy for the latter. For single transformations or actions of $\mathbf{Z}^{d}$, differential Kolmogorov–Sinai entropy for stationary $\mathbf{R}$-valued processes was studied in [78], prompted by earlier work in information theory such as [119]. Section 6 of [78] includes some exact calculations for Gaussian systems and linear transformations between them which boil down to applications of Theorem S itself. However, this differential version of Kolmogorov–Sinai entropy starts to behave quite wildly beyond those Gaussian examples, and for those examples alone we might as well stay within the setting of representation theory. Overall, an analogy at the level of intuition seems more revealing than either the Koopman or the Gaussian construction for our work in this paper.
Beyond Theorem A, we consider two other notions of entropy from ergodic theory, and develop their analogs for unitary representations of groups or representations of other C*-algebras. The first of these notions is Kieffer’s definition of entropy using a ‘random past’ from [96, Theorem 3], which leads to Theorem B. The second is Bowen’s ‘sofic entropy’ from [25], which leads to Theorem C. In both cases, we show that these notions of entropy are given by $\log\Delta\varphi_{\mathrm{ac}}$ for some positive functional $\varphi$, except in Theorem C a certain degeneracy may occur and then the entropy equals $\infty$- ∞ (also reflecting a known feature of sofic entropy).
Each of these theorems may be seen as a different ‘non-commutative’ version of Theorem S. In ergodic theory, various predecessors of our results are already discussed this way in the literature, for example in [110, 104, 80].
@@ 1.3 Subdiagonal subalgebras and random orders
where $P_{n}$ is the orthogonal projection from $L^{2}(\mu)$ to $\mathrm{span}\{z,\dots,z^{n}\}$ [136, Theorem 1.5.11]. It follows that
@@
This orthogonal projection is the closest point of $N$ to the function $1_{\mathbf{T}}$, so Theorem S is equivalent to
@@
This is the form in which Szegő’s theorem most often appears in the literature on analytic functions, such as in [67, Section V.8] or [68, Theorem IV.3.1]. Those references give essentially Szegő’s own proof, which he published first for the special case $\mu\ll m$, and then much later for the general case by incorporating arguments of Kolmogorov and Krein to handle $\mu_{\mathrm{sing}}$. Two further variations on this proof are recounted in [136, Sections 2.4 and 2.5]. Meanwhile, Verblunsky gave the first proof of the general case of Theorem S in [152, 151] (see also [136, Section 2.3]).
@@
See, for instance, [157, Theorem 4.14].
To generalize Szegő’s theorem to positive definite functions on $\mathbf{Z}^{d}$, one can use the ‘past’ defined by a lexicographic ordering. This approach was developed by Helson and Lowdenslager in [85, 86] (with some later refinements in [109]). See also [105, 51] for alternative proofs under some additional assumptions on the positive definite functions. In ergodic theory, the same use of the lexicographic ordering appeared in some early works on the entropy of measure-preserving $\mathbf{Z}^{d}$-actions such as [39, 91].
This idea has a further natural generalization to any countable group $\Gamma$ that admits a left-invariant total order. But some countable groups do not, including all groups that are not torsion-free. To remove the need for this assumption, one can instead couple to a stationary random ordering of the group. All countable groups admit at least one of these: the ‘Bernoulli random order’, described in Example 7.5 below.
In ergodic theory, conditioning on the ‘past’ of a stationary random order can sometimes serve as a replacement for the formula (1.5). If $\Gamma$ is amenable, this idea gives rise to another classic formula for the entropy rate of a process due to Kieffer [96]. Our next main theorem develops the analog of this idea for positive definite functions. We find once again that the resulting quantity always agrees with the expected Fuglede–Kadison determinant, even if $\Gamma$ is not amenable.
Before formulating the theorem precisely, let us motivate it further via a related finite-dimensional calculation. Let $n$ and $k$ be positive integers. Consider $nk$ vectors in some Hilbert space $H$ indexed as an $n$-by-$k$ array, say
@@
Formula (1.6) remains valid under any re-ordering of the tuples $V_{1}$, …, $V_{n}$, and so we can now take an expectation over a uniform random order on the right-hand side. Then the symmetry of the random order lets us replace the average of $n$ terms with a single average over orders. Using $\omega$ to denote a permutation of $1$, …, $n$, and now writing $R_{\omega}$ for the orthogonal projection onto the random subspace
@@
From the viewpoint of ergodic theory, Theorem B is somewhat surprising, because for a non-amenable group $\Gamma$ it gives much better behaviour than the original notion of random-order entropy for a finite-valued stationary processes. Random-order entropy does agree with Rokhlin or sofic entropy (discussed below) for special examples of processes, such as those in [12, 4]. But in general it is only an upper bound for those quantities, and it need not be invariant under isomorphism [133, Section 7]. However, for a positive definite function $\varphi$ on a group $\Gamma$ for which both Theorems B above and C below can be applied, they always give the same value $\log\Delta\varphi_{\mathrm{ac}}$ for their respective notions of entropy. Since Theorem B concerns the analog of random-order entropy, and Theorem C concerns the analog of sofic entropy, this suggests that the worst pathologies from ergodic theory do not appear for unitary representations.
It turns out that most of the work towards Theorem B can already be found in the literature. It takes place in the context of Arveson’s theory of ‘subdiagonal subalgebras’ of finite von Neumann algebras [9]. These are a noncommutative generalization of the inclusion of the Hardy space $H^{\infty}(\mathbf{T})$ in $L^{\infty}(\mathbf{T})$. As such, they provide an abstract notion of a ‘past’ in a general finite von Neumann algebra. We recall their definition precisely in Subsection 7.5.
Arveson conjectured a generalization of Szegő’s theorem to this noncommutative setting in [9, item 4.4($\gamma$)]. As far as I know, this was the first reference to suggest that the log-integral should be replaced with a Fuglede–Kadison determinant. Arveson proved this generalization in a few of his examples in [9, Section 5], including some that recover Helson and Lowdenslager’s results for $\mathbf{Z}^{d}$. He also showed how to construct subdiagonal subalgebras for any countable group with an invariant total order, but did not prove his Szegő formula in this generality.
Arveson’s conjectured Szegő theorem was finally proved in general by Labuschagne in [100], in the equivalent form of an abstract ‘Jensen formula’. We recall this theorem as Theorem 7.3 below.